OpenStudy (caozeyuan):

show that the area under r=a*sin(3theta) is 1/12 * a*pi, where theta is between 0 and 1/3 pi, inclusively

OpenStudy (caozeyuan):

sorry, it should be 1/12*a^2*pi

OpenStudy (caozeyuan):

@amistre64 help!!!!

OpenStudy (amistre64):

\[\int^{\beta}_{\alpha}\frac {[r(t)]^2}2~dt\]

OpenStudy (amistre64):

just remember the formula for the area of a circle:\[\frac12~2pi~r^2\]2pi represent the radians traveled, so replace 2pi with dt

OpenStudy (amistre64):

do you recall the integration of sin^2(x) ?

OpenStudy (caozeyuan):

nope, not at all

OpenStudy (amistre64):

well, last time we tangoed ... you said you hadnt learnded by parts yet. Have you learnded it yet?

OpenStudy (caozeyuan):

ye, I learn that but I don't know to use it on the sin^2 x function

OpenStudy (amistre64):

\[s^2(3t)=s(3t)~s(3t)\] u = s(3t) v = -c(3t) du = 3 c(3t) dv = s(3t) int s^2(3t) = -s(3t) c(3t) - int -3 c(3t) c(3t) dt

OpenStudy (amistre64):

define your parts and run with it ... youll see that by the second run thru you are going in circles ..... but theres a way out of it

OpenStudy (amistre64):

\[\int s^2(3t)~dt=-s(3t)~c(3t)/3-\int-3c(3t)~c(3t)/3~dt\] \[\int s^2(3t)~dt=-s(3t)~c(3t)/3+\int c(3t)~c(3t)~dt\] run thru by parts with the cosines now u = c(3t) v = s(3t)/3 du = -3 s(3t) dv = c(3t) \[\int s^2(3t)~dt=-s(3t)~c(3t)/3+\left(s(3t)~c(3t)/3-\int -3s(3t)~s(3t)/3~dt\right)\] \[\int s^2(3t)~dt=-s(3t)~c(3t)/3+s(3t)~c(3t)/3+\int s^2(3t)~dt\] pretty sure ive got an error in that someplace .... coding and mathing tend to mess things up for me :) we should have a multiple of the original integral on the end so that we can subtract it from both sides

OpenStudy (amistre64):

might have to work up some trig identity stuff .... sin^2(3t) = 1-cos^2(3t) cos(2x) = cos^2(x) - sin^2(x) = cos^2(x) - (1 - cos^2(x)) = 2cos^2(x) - 1 (cos(2x) + 1)/2 = cos^2(x) , let x = 3t -cos^2(3t) = -(cos(6t)+1)/2 \[\int sin^2(3t)dt=\int 1-\frac12cos(6t)~dt=t-\frac12~\frac16sin(6t)\]

OpenStudy (amistre64):

so, this amounts to\[\frac12a^2\left( t-\frac1{12}sin(6t)-[t-\frac1{12}sin(6t)]\right)^{pi/3}_{0}\] \[\frac12a^2\left( \frac{pi}3-\frac1{12}sin(6*\frac{pi}3)-[0-\frac1{12}sin(6.0)]\right)\] \[\frac12a^2\left( \frac{pi}3-\frac1{12}(0)\right)\]

OpenStudy (amistre64):

i missed the +1 when i coded up the bottom of the other post :/

OpenStudy (amistre64):

\[\int sin^2(3t)dt=\int 1-\frac12(cos(6t)+1)~dt=t-\frac12~\frac16sin(6t)-\frac12t\] \[\frac12t-\frac1{12}sin(6t)\]